The maximum safe current that can pass through a resistor of 2 kΩ rated at 0.25 Watts is

1
Expert's answer
2021-12-01T22:46:12-0500

Solution;

"P=I^2R"

Given;

"R=2000\Omega"

"P=0.25"

Therefore;

"I^2=\frac PR"

"I=\sqrt{\frac PR}=\sqrt{\frac{0.25}{2000}}=12.5mA"